OpenStudy (anonymous):

What is the solution of the system of equations? y = –2x + 5 y = –2x + 20

OpenStudy (anonymous):

no solution (1, 3) (0, 20) infinitely many solutions

pooja195 (pooja195):

Use substituion we already know the value of y right? Plug E1 into E2 -2x+5= –2x + 20 solve for x

OpenStudy (anonymous):

so is it like this -2(2)+5 = _2(2)+20

pooja195 (pooja195):

Well you cant solve the equation :) Add 2x to both sides. −2x+5+2x=−2x+20+2x 5=20 Subtract 5 from both sides. 5−5=20−5 0=15

OpenStudy (anonymous):

sorry but I don't get it

pooja195 (pooja195):

Ok when you have m=m <---meaning 2 of the SAME numbers equaling each other it means that there are infinite solutions m=1 <-----one solutions m=n <---No solutions

pooja195 (pooja195):

in this case once we solve \[\huge~\rm~-2x+5= –2x + 20 \] We get 0=15 therefore there are no solutions to this system of equations

OpenStudy (anonymous):

oh okay

pooja195 (pooja195):

Do you understand?

OpenStudy (anonymous):

yes I understood a bit, thanks

pooja195 (pooja195):

Let me know if anything is unclear :) And your welcome :)